OpenStudy (selena2345):

If a trip between work and home takes 15 minutes each way, how much time will be spent in a 5 day work week traveling back and forth to work ? Justify your reasoning

OpenStudy (selena2345):

@tkhunny can you please help me ?

OpenStudy (tkhunny):

How much travel time for one day?

OpenStudy (selena2345):

i don't know @tkhunny

OpenStudy (tkhunny):

That is an inadequate response. You must study it in your mind and find a reasonable answer. 15 minutes each way is how much in total?

OpenStudy (darkbluechocobo):

15+15=30

OpenStudy (darkbluechocobo):

thats 1 day

OpenStudy (darkbluechocobo):

2 days is 60 3 days is 90 4 days is 120 5 days 150

OpenStudy (selena2345):

ok @DarkBlueChocobo

OpenStudy (darkbluechocobo):

150 is the answer because its 30 minutes back and forth each day in total

OpenStudy (selena2345):

how @DarkBlueChocobo

OpenStudy (darkbluechocobo):

what do you mean?

OpenStudy (kendricklamar2014):

150 minutes

OpenStudy (kendricklamar2014):

30*5= 150 minutes

OpenStudy (darkbluechocobo):

you add 30 each day because it takes 30 minutes to go to work and come back

OpenStudy (kendricklamar2014):

Exactly ^

OpenStudy (selena2345):

can you show me how?

OpenStudy (kendricklamar2014):

30 * 5= 150 30 = How much time it takes to get to work and back to his house. 5= 5 days

OpenStudy (selena2345):

what grade are you in ? @KendrickLamar2014

OpenStudy (kendricklamar2014):

Im a Freshman, so 9th

OpenStudy (selena2345):

me too

OpenStudy (selena2345):

thank you so much @KendrickLamar2014

OpenStudy (kendricklamar2014):

No problem